# SAR-TEXT: A Large-Scale SAR Image-Text Dataset Built with SAR-Narrator and Progressive Transfer Learning

### v1

@@ 4.3.4 SAR-RS-CoCa Caption Generation Qualitative Assessment
In summary, SAR-RS-CoCa performs well in both types of tasks: for scene understanding tasks, it can generate descriptions with global semantics; for object recognition tasks, it can provide accurate and practical textual information. This demonstrates the effectiveness of the SAR-TEXT dataset in supporting generative model training, while also validating the applicability and robustness of the incremental transfer learning strategy in the field of SAR imagery.
4.4 Ablation Study
Table 6: Comparison of Two Training Strategies on the HRSCD Test Set
@@ 4.4 Ablation Study
  • Single-stage mixed training (Single-SAR-RS-CLIP): This training scheme integrates 210,000 optical remote sensing image–text pairs from the HQRS-IT-210K dataset and 130,000 SAR image–text pairs from the SAR-TEXT dataset to jointly train the RS-CLIP model. All samples are uniformly fed into the model, with the data loader randomly selecting instances from the combined dataset in each training batch. The resulting model is referred to as Single-SAR-RS-CLIP.
  • Two-stage progressive training (SAR-RS-CLIP):This training scheme first pre-trains the CLIP model on the HQRS-IT-210K dataset to adapt it to the optical remote sensing image–text matching task. The model is then further fine-tuned using the SAR-TEXT dataset to complete the transfer to SAR image–text matching. The resulting model is referred to as SAR-RS-CLIP.
Tables 6 and 7 show the performance comparison of the two training strategies on the HRSCD and OSdataset_512 test sets. The evaluation metrics cover the recall rates R@1, R@5, and R@10 for both image retrieval to text (i2t) and text retrieval to image (t2i), as well as the overall mean recall.
On the HRSCD test set, SAR-RS-CLIP outperforms Single-SAR-RS-CLIP across all metrics. In the i2t direction, all three recall metrics show improvements, with a relative increase of +3.93% in i2t-R@10. Improvements in the t2i direction are equally notable, with the t2i-R@10 metric improving by +3.56%. The overall average recall rate increased from 9.44% to 11.52%, representing a relative increase of 22.0%, indicating that the two-stage training strategy demonstrates superior semantic alignment and retrieval capabilities in medium-to-low resolution remote sensing scenarios.
@@ 4.5.2 Qualitative Analysis of SAR-GPT of SAR-GPT
Figure 12: Example of Human–Machine Dialogue with SAR-GPT.
Overall, both quantitative and qualitative evaluations demonstrate that SAR-GPT exhibits strong image analysis and natural language generation capabilities, confirming its effectiveness in cross-modal alignment and SAR image understanding. These results further underscore the practicality of the SAR-TEXT dataset for downstream tasks.

### v3

@@ 4.3.4 SAR-RS-CoCa Caption Generation Qualitative Assessment
Table 6: Comparison of Two Training Strategies on the HRSID Test Set
@@
In summary, SAR-RS-CoCa demonstrates strong capabilities in SAR image captioning. For natural scene imagery, it can generate coherent descriptions that capture key global semantic elements such as land cover types and structural patterns. For images containing specific targets, the model can produce textual descriptions that include not only object categories but also indicative positional or quantitative information, thus enriching the interpretability of the captions. These results highlight the effectiveness of the SAR-TEXT dataset in supporting caption generation for SAR images, and also suggest that the progressive transfer learning strategy provides a promising approach for enhancing the semantic expressiveness of generative models in SAR-related applications.
4.4 Ablation Study
@@
  • Single-stage mixed training (Single-SAR-RS-CLIP): This training scheme integrates 210,000 optical remote sensing image–text pairs from the HQRS-IT-210K [9] dataset and 130,000 SAR image–text pairs from the SAR-TEXT dataset to jointly train the RS-CLIP model. All samples are uniformly fed into the model, with the data loader randomly selecting instances from the combined dataset in each training batch. The resulting model is referred to as Single-SAR-RS-CLIP.
  • Two-stage progressive training (SAR-RS-CLIP):This training scheme first continual pretraining the CLIP model on the HQRS-IT-210K dataset to adapt it to the optical remote sensing image–text matching task. The model is then further fine-tuned using the SAR-TEXT dataset to complete the transfer to SAR image–text matching. The resulting model is referred to as SAR-RS-CLIP.
Tables 6 and 7 show the performance comparison of the two training strategies on the HRSID and OSdataset_512 test sets. The evaluation metrics cover the recall rates R@1, R@5, and R@10 for both image retrieval to text (i2t) and text retrieval to image (t2i), as well as the overall mean recall.
On the HRSID test set, SAR-RS-CLIP outperforms Single-SAR-RS-CLIP across all metrics. In the i2t direction, all three recall metrics show improvements, with a relative increase of +3.93% in i2t-R@10. Improvements in the t2i direction are equally notable, with the t2i-R@10 metric improving by +3.56%. The overall average recall rate increased from 9.44% to 11.52%, representing a relative increase of 22.0%, indicating that the two-stage training strategy demonstrates superior semantic alignment and retrieval capabilities in medium-to-low resolution remote sensing scenarios.
@@ 4.5.1 Quantitative Evaluation of SAR-GPT
Figure 12: Example of Human–Machine Dialogue with SAR-GPT.
@@ 4.5.2 Qualitative Analysis of SAR-GPT of SAR-GPT
Overall, our preliminary quantitative and qualitative observations suggest that SAR-GPT can leverage the SAR-TEXT dataset to produce reasonable image–text responses, indicating the potential of SAR-TEXT to support GPT-style multimodal models in the SAR domain. However, the current results should be interpreted as exploratory rather than conclusive: The SAR-GPT training process remains relatively simple, the evaluation benchmarks are limited in scale and coverage, and the generated answers sometimes show stylistic bias or omit fine-grained details. We regard this study as an initial attempt to probe the practicality of SAR-TEXT for large language model adaptation, while leaving a more systematic exploration of SAR-GPT training and evaluation to future work.